# Quasinormal modes and bound states of massive scalar fields in generalized Bronnikov-Ellis wormholes

### 05 Mar 2024 v2

@@ Appendix A Numerical values
In this appendix we provide some numerical values of the QNFs via CFM for the generalized Bronnikov-Ellis wormholes. In Table 2 and 3 we show the QNFs for the fundamental mode for $M/b_{0}=0.1$, and $M/b_{0}=0.25$, respectively, and in Table 4 and 5 we show the QNFs for the first overtone ($n=1$) for $M/b_{0}=0.1$, and $M/b_{0}=0.25$, respectively.
Table 2: Fundamental $n=0$ QNFs $\omega b_{0}$ for scalar perturbations for the generalized Bronnikov-Ellis wormhole with $M/b_{0}=0.1$, for several values of the angular momentum and the mass of the scalar field using the CFM.
@@
Table 3: Fundamental $n=0$ QNFs $\omega b_{0}$ for scalar perturbations for the generalized Bronnikov-Ellis wormhole with $M/b_{0}=0.25$, for several values of the angular momentum and the mass of the scalar field using the CFM.
@@
Table 4: First overtone $n=1$ QNFs $\omega b_{0}$ for scalar perturbations for the generalized Bronnikov-Ellis wormhole with $M/b_{0}=0.1$, for several values of the angular momentum and the mass of the scalar field using the CFM.
@@
Table 5: First overtone $n=1$ QNFs $\omega b_{0}$ for scalar perturbations for the generalized Bronnikov-Ellis wormhole with $M/b_{0}=0.25$, for several values of the angular momentum and the mass of the scalar field using the CFM.
@@ Appendix B Comparison of CFM and WKB method
Table 6: QNFs $\omega b_{0}$ for massive scalar field perturbations $m=0.5$ for the generalized Bronnikov-Ellis wormhole with $M/b_{0}=0.1$, $n=0,1$, for several values of the angular momentum $\ell$ of the scalar field using the CFM and the six-order WKB approximation.

### 10 Apr 2024 v3

@@ Appendix A Numerical values
In this appendix we provide some numerical values of the QNFs via CFM for the wormhole spacetime described by Eq. (13). In Table 4 and 5 we show the QNFs for the fundamental mode for $M/b_{0}=0.1$, and $M/b_{0}=0.25$, respectively, and in Table 7 and 8 we show the QNFs for the first overtone ($n=1$) for $M/b_{0}=0.1$, and $M/b_{0}=0.25$, respectively.
Table 3: Fundamental $n=0$ QNFs $\omega b_{0}$ for scalar perturbations for the Bronnikov-Ellis wormhole for several values of the angular momentum and the mass of the scalar field using the CFM.
$\ell$ $mb_{0}=0$ $mb_{0}=0.2$ $mb_{0}=0.4$ $mb_{0}=0.6$
0 0.68136719-0.61775345 i 0.69762539-0.60335666 i 0.74791262-0.56278891 i 0.83479189-0.50421780 i
1 1.57270839-0.52970088 i 1.58410509-0.52589000 i 1.61799977-0.51487339 i 1.67350804-0.49779565 i
2 2.54665642-0.51266881 i 2.55419456-0.51115579 i 2.57669677-0.50669188 i 2.61383299-0.49949302 i
3 3.53428741-0.50692165 i 3.53982819-0.50612818 i 3.55640289-0.50376936 i 3.58387030-0.49990838 i
5 5.52230920-0.50295823 i 5.52589995-0.50263141 i 5.53665868-0.50165471 i 5.55454509-0.50003931 i
10 10.51183964-0.50083853 i 10.51373778-0.50074811 i 10.51943015-0.50047714 i 10.52891067-0.50002650 i
15 15.50804389-0.50038767 i 15.50933215-0.50034610 i 15.51319628-0.50022147 i 15.51963438-0.50001396 i
$\ell$ $mb_{0}=0.8$ $mb_{0}=1.0$ $mb_{0}=1.2$ $mb_{0}=1.4$
0 0.95710482-0.43978143 i 1.10771470-0.37998677 i 1.27720491-0.32956100 i 1.45807345-0.28868020 i
1 1.74918582-0.47625873 i 1.84312741-0.45198450 i 1.95313933-0.42652616 i 2.07694738-0.40110069 i
2 2.66507432-0.48988928 i 2.72972309-0.47828709 i 2.80695095-0.46512795 i 2.89584111-0.45085047 i
3 3.62200113-0.49464557 i 3.67048645-0.48811154 i 3.72894874-0.48045895 i 3.79695454-0.47185363 i
5 5.57949282-0.49780347 i 5.61141062-0.49497195 i 5.65018391-0.49157531 i 5.69567673-0.48764897 i
10 10.54216920-0.49939763 i 10.55919163-0.49859255 i 10.57995997-0.49761382 i 10.60445239-0.49646451 i
15 15.52864326-0.49972388 i 15.54021845-0.49935166 i 15.55435426-0.49889785 i 15.57104374-0.49836312 i
Table 4: Fundamental $n=0$ QNFs $\omega b_{0}$ for scalar perturbations for the wormhole spacetime described by Eq. (13) with $M/b_{0}=0.1$, for several values of the angular momentum and the mass of the scalar field using the CFM.
@@
Table 5: Fundamental $n=0$ QNFs $\omega b_{0}$ for scalar perturbations for the wormhole spacetime described by Eq. (13) with $M/b_{0}=0.25$, for several values of the angular momentum and the mass of the scalar field using the CFM.
@@
Table 6: First overtone $n=1$ QNFs $\omega b_{0}$ for scalar perturbations for the Bronnikov-Ellis wormhole for several values of the angular momentum and the mass of the scalar field using the CFM.
$\ell$ $mb_{0}=0$ $mb_{0}=0.2$ $mb_{0}=0.4$ $mb_{0}=0.6$
0 0.46716556-2.17646191 i 0.46906191-2.16766276 i 0.47488384-2.14108791 i 0.48505482-2.09619201 i
1 1.25582759-1.70245148 i 1.26145961-1.69485057 i 1.27859466-1.67213708 i 1.30794582-1.63461322 i
2 2.34497715-1.57249831 i 2.35086218-1.56856180 i 2.36853730-1.55685647 i 2.39805487-1.53769317 i
3 3.39009976-1.53721055 i 3.39499204-1.53499538 i 3.40965523-1.52839415 i 3.43404739-1.51753791 i
5 5.43090081-1.51524731 i 5.43431671-1.51429486 i 5.44455525-1.51144721 i 5.46158899-1.50673327 i
10 10.46413347-1.50422829 i 10.46600591-1.50395918 i 10.47162139-1.50315267 i 10.48097437-1.50181129 i
15 15.47575755-1.50194589 i 15.47703778-1.50182166 i 15.48087787-1.50144912 i 15.48727599-1.50082884 i
$\ell$ $mb_{0}=0.8$ $mb_{0}=1.0$ $mb_{0}=1.2$ $mb_{0}=1.4$
0 0.50037244-2.03202248 i 0.52218113-1.94715585 i 0.55270365-1.83962607 i 0.59567461-1.70691857 i
1 1.35067753-1.58289857 i 1.40832928-1.51810061 i 1.48262974-1.44202256 i 1.57514763-1.35732390 i
2 2.43947742-1.51158301 i 2.49284354-1.47922345 i 2.55812690-1.44147368 i 2.63519525-1.39931666 i
3 3.46809539-1.50263950 i 3.51169079-1.48398519 i 3.56468519-1.46192352 i 3.62688609-1.43685161 i
5 5.48537240-1.50020039 i 5.51584212-1.49191323 i 5.55291737-1.48195215 i 5.59650057-1.47041133 i
10 10.49405566-1.49993922 i 10.51085246-1.49754225 i 10.53134838-1.49462776 i 10.55552357-1.49120463 i
15 15.49622907-1.49996173 i 15.50773287-1.49884904 i 15.52178191-1.49749240 i 15.53836956-1.49589379 i
Table 7: First overtone $n=1$ QNFs $\omega b_{0}$ for scalar perturbations for the wormhole spacetime described by Eq. (13) with $M/b_{0}=0.1$, for several values of the angular momentum and the mass of the scalar field using the CFM.
@@
Table 8: First overtone $n=1$ QNFs $\omega b_{0}$ for scalar perturbations for the wormhole spacetime described by Eq. (13) with $M/b_{0}=0.25$, for several values of the angular momentum and the mass of the scalar field using the CFM.
@@ Appendix B Comparison of CFM and WKB method
Table 9: QNFs $\omega b_{0}$ for massive scalar field perturbations $m=0.5$ for the wormhole spacetime described by Eq. (13) with $M/b_{0}=0.1$, $n=0,1$, for several values of the angular momentum $\ell$ of the scalar field using the CFM and the six-order WKB approximation.